OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 11 -3x + 2y = 1 a. 2 b. 4 c. y can be any number as there are infinitely many solutions to the system. d. There is no y value as there is no solution to the system.

Vocaloid (vocaloid):

add the two equations together

OpenStudy (anonymous):

didn't we do this one already?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

okay well chose one of the equations and set it equal to x

OpenStudy (anonymous):

3x+11= y?

OpenStudy (anonymous):

yea sure either one so on that one just subtract 11 and divide by 3

OpenStudy (anonymous):

then you'll get (y-11)/3 = x

OpenStudy (anonymous):

and now plug the (y-11)/3 into the other equation as if it was x

OpenStudy (anonymous):

So (-3x +y -11)/3 + 2y=1 ?

OpenStudy (anonymous):

@Vocaloid